OpenStudy (anonymous):

GIVING MEDAL PLEASE HELP ME i'm trying to make a study guide so please help! i'll give u a medal

OpenStudy (anonymous):



OpenStudy (anonymous):

Its for my upcoming test so i need to finish my study guide.

OpenStudy (shinalcantara):

Substitute the values of x and y \[\frac{ x }{ y }=\frac{ 8 }{ 6 }\] Reduce it to lowest form. The largest common factor both numbers have is 2 divide both the numerator and denominator by 2 \[\frac{ 8 \div 2 }{ 6 \div 2 } = ?\]

OpenStudy (ahsome):

Your equation is: \[\frac{x}{y}\] Where: \(x=8\) \(y=6\) Firstly, put those numbers into the equation \[\frac{x}{y}\]\[=\frac{8}{6}\]Now, what is the biggest common factor between the two numbers? \(2\). So simply divide both numbers by 2 \[=\frac{8}{6}\]Simplify\[=\frac{8\div2}{6\div2}\]\[=\frac{4}{3}\] The anser is \(\color{blue}{\text{B}}\), \(\color{red}{\frac{4}{3}}\)

OpenStudy (ahsome):

\(\color{green}{\text{Answer*}}\)

OpenStudy (anonymous):

yes thx

OpenStudy (ahsome):

No problem @TaigaA :)